OpenStudy (anonymous):

If A lies in third quadrant and 3tanA -4 =0 then 5sin2A + 3sinA +4cosA is equal to

OpenStudy (loser66):

3tan a -4=0, tan a = 4/3, A is in third quadrant, then, cos and sin have negative value. (let it there)

OpenStudy (anonymous):

ok

OpenStudy (loser66):

the hypotenuse of the triangle is 5, right?

OpenStudy (anonymous):

i got it

OpenStudy (loser66):

then , sin2A = 2 sin A cos A

OpenStudy (anonymous):

answer is 0 yupiie

OpenStudy (anonymous):

i got it tititi

OpenStudy (anonymous):

thanks

OpenStudy (loser66):

so that sin A =\(\dfrac{3}{5}\)

OpenStudy (loser66):

ok